TheJag18:

A consumer owes $2,550 on a credit card with a 21.9% annual interest rate. Assuming the consumer makes fixed payments and does not charge any more purchases with the card, what is the total monthly payment needed to pay off the card in three years? Show all work and round your answer to the nearest penny.

Phantomdex:

I believe your answer would be $97.25/monthly. This is because it gives us both what the consumer owes, and the annual interest rate.

TheJag18:

can you show the full work? This needs all of it

Phantomdex:

I think I miscalculated anyways so hold on

TheJag18:

kk

Phantomdex:

Given: P = 2,550 R = 21.9% and T = 3 years I =PTR/100 = 2559 x 21.9 x 3/100 = 1675.35 Amount = P + I = $2550 + 1675.35 =$4225.35 Monthly payment = amount/3 x 12 months = 4225.35/3 x 12 =4225.35/36 = $117.32 = $117.40 (nearest penny)

TheJag18:

Awsome, thank you :)

Phantomdex:

You're welcome